# Microphysical Regulation of Non-Ideal MHD in Weakly-Ionized Systems: Does the Hall Effect Matter?

### v1

@@ 2.1 Derivation
Figure 1: Illustration of the regimes in density $\bar{n}=\bar{\rho}/m_{p}$ and magnetic field strength $\bar{B}\equiv|\bar{\bf B}|$, in a weakly-ionized gas ($n_{\pm}\ll n$), where different non-ideal MHD terms derived in § 2.1 (Ohmic resistivity $\eta_{O}{\bf J}$, Hall $\eta_{H}{\bf J}\times{\bf b}$, and ambipolar drift $\eta_{A}{\bf J}\times{\bf b}\times{\bf b}$) would each be relatively more important in the induction equation. Quantities like temperature and ionization fractions at each $B$, $n$ are estimated with a simple 5-species toy model (§ 2.4.3). Left: The case with “classical” non-ideal coefficients (§ 2.2, Eq. 8), which are valid only if the drift velocity between charge carriers $v_{\rm drift}\sim|{\bf J}|/n_{-}|q_{-}|\sim cB/4\pi n_{-}|q_{-}|\ell_{B}$ ($\ell_{B}\sim|{\bf B}|/|\nabla{\bf B}|$) and “slip” velocity between neutrals and carriers (Eq. 9) are much less than the thermal velocities $v_{T}$. This in turn requires that the magnetic gradient scale length $\ell_{B}$ is much larger than some $\ell_{\rm crit}$ (Eq. 13). Right: The case with corrected “effective” coefficients (§ 2.3; Eqs. 15-16), which account for superthermal drift and its effect both on direct particle collision rates and effective collisions via scattering (anomalous resistivity), assuming drift velocities are superthermal (here $v_{\rm drift}=10\,v_{T}$ for illustration), or $\ell_{B}\lesssim\ell_{\rm crit}$. When the drift becomes superthermal, the effective Ohmic term is always larger than the Hall term, and increases greatly in relative importance even at low densities. Where it predominates, we label whether the anomalous ($\omega_{\rm an}$) or direct collision ($\omega_{\rm coll}$) term is most important.
We begin with a derivation of the non-ideal MHD equations, following e.g. Cowling (1976); Ichimaru (1978); Nakano & Umebayashi (1986). We follow the usual kinetic approach by computing moments of the distribution function of each species, assuming infinitesimally small averaging volumes (but implicitly larger than the interparticle separation and Debye length), with bulk velocities $\ll c$, quasi-neutrality, and collisional/exchange reactions being mass and momentum conserving (e.g. rest mass loss via photons is negligible). For an arbitrary set of species $j$ (which should formally include all isotopes/masses/ionization states/charges) this gives,
@@
Now consider a three-component system with a dominant set of neutrals, positive and negative charge carriers (labeled $n$, $+$, and --, respectively). We can repeat this derivation for an arbitrary set of species, but all of the salient dimensional scalings and intuition will be identical so the expressions are much more straightforward in this limit. Define an “ambipolar current” ${\bf J}_{d}\equiv-(q_{-}\rho_{c}/m_{-})\,({\bf u}_{+}-{\bf u}_{n})$, where $\rho_{c}\equiv\rho_{+}+\rho_{-}$. We can then express the three velocities ${\bf u}_{n,+,-}$ in terms of ${\bf U}$, ${\bf J}$, and ${\bf J}_{d}$: ${\bf u}_{n}={\bf U}-(m_{-}/q_{-}\rho)\,({\bf J}-{\bf J}_{d})$, ${\bf u}_{+}={\bf U}-(m_{-}/q_{-}\rho)\,[{\bf J}+(\rho_{n}/\rho_{c}){\bf J}_{d}]$, ${\bf u}_{-}={\bf U}-(m_{-}/q_{-}\rho)\,[-\{(\rho_{+}+\rho_{n})/\rho_{-}\}{\bf J}+(\rho_{n}/\rho_{c}){\bf J}_{d}]$ (Cowling, 1976; Pinto et al., 2008), and rewrite our three Eq. 1 for the evolution of ${\bf u}_{n,+,-}$ with an equivalent set for ${\bf U}$, ${\bf J}$, ${\bf J}_{d}$ via substitution. With Eq. 2 for ${\bf U}$, we have:
@@
${\bf G}_{0}$ $\equiv\frac{|q_{-}|}{m_{-}}[{\xi_{n}\nabla\cdot\boldsymbol{\Pi}-\nabla\cdot\boldsymbol{\Pi}_{n}+\xi_{n}({\bf a}_{e,-}\rho_{-}+{\bf a}_{e,+}\rho_{+}-{\bf a}_{e,n}\rho_{c})}]$
@@
$\quad+{{\bf D}_{t0}}+{\bf G}_{2}-\frac{|q_{-}|}{m_{-}}\nabla\cdot(\boldsymbol{\Pi}_{-}+\epsilon\boldsymbol{\Pi}_{+}){\Bigr{]}}\ ,$
@@
Note that $\omega_{O}$ is similar to an effective collision frequency $\omega^{\rm coll}$; $\omega_{H}$ to the gyrofrequency $\sim\Omega_{-}$, and $\omega_{A}$ scales as $\sim\Omega_{-}\Omega_{+}/\omega^{\rm coll}$.
@@
$\partial_{t}\bar{\bf B}$ $=-c\left[\nabla\times\langle{\bf E}\rangle_{\mathcal{V}}+\langle\nabla\times\delta{\bf E}\rangle_{\mathcal{V}}\right]$ (7)
=×$=\nabla\times\left[\bar{\bf U}\times\bar{\bf B}+\langle\delta{\bf U}\times\delta{\bf B}\rangle_{\mathcal{V}}+\bar{a}_{O}\bar{\bf J}+\langle\delta a_{O}\delta{\bf J}\rangle_{\mathcal{V}}+...\right]$
+×$+\langle\nabla\times\left[\delta{\bf U}\times\bar{\bf B}+\bar{\bf U}\times\delta{\bf B}+...\right]\rangle_{\mathcal{V}}$
where the “…” represents the appropriate (and extensive) expansion of Eq. 5 for all fluctuating terms. Note that, as in large-scale dynamo theory (Rincon, 2019), this is not a series expansion, so we cannot prima facia assume the $\delta$ terms are small and drop them. Also of importance is that, given some state vector $\boldsymbol{\Psi}\equiv({\bf U},{\bf B},{\bf J},{\bf J}_{d},...)$, variables like $\omega_{i}$ or $a_{i}$ will not satisfy $\bar{a}_{i}=a_{i}(\bar{\boldsymbol{\Psi}})$, because they are non-linear functions of the state variables. Evaluating terms like $\langle\delta{\bf U}\times\delta{\bf B}\rangle_{\mathcal{V}}$ requires integrals over the entire spatial and time spectrum of fluctuations knowing the appropriate $N$-th order correlation functions between the fluctuations, so we stress that Eq. 7 is no more integrable than Eq. 5, unless we impose additional assumptions.
@@ 2.2 The “Standard” Non-Ideal MHD Approximations
To arrive at the standard formulation of one/two-fluid non-ideal MHD, we simplify Eq. 7 by making the following assumptions (see Ichimaru, 1978; Norman & Heyvaerts, 1985; Nakano & Umebayashi, 1986; Wardle & Ng, 1999; Tassis & Mouschovias, 2005, 2007; Kunz & Mouschovias, 2009). (1) Assume (given dimensional arguments since velocities are non-relativistic: $|{\bf U}|,\,c_{s},\,v_{\rm drift},\,v_{A}\ll c$) that displacement currents can be neglected so ${\bf J}\rightarrow{\bf J}_{A}\equiv(c/4\pi)\nabla\times{\bf B}$. (2) Drop all terms from the second-line of Eq. 5. (a) Assume temporal/advection frequencies $D_{t}$ are much smaller than the collision+gyro frequencies $\omega_{i}$ ($|D_{t}|\ll\omega_{i}$), usually justified by assuming $|D_{t}|\sim 1/\Delta t_{\rm macro}$ varies only on “macroscopic” timescales because all currents come into local steady-state ($D_{t}\bar{\bf X}\rightarrow 0$) rapidly. (b) Assume that the battery terms are negligible, i.e. external accelerations are negligible compared to the typical non-ideal terms, $|{\bf a}_{\rm ext}|\ll|\omega_{i}v_{\rm drift}|$, ionization fractions are small, and the difference between external acceleration/pressure forces on neutrals and charge carriers is small. (c) Assume the stresses are similarly negligible. Noting $\nabla\cdot(\boldsymbol{\Pi}_{-}+\epsilon\boldsymbol{\Pi}_{+})\sim\rho_{-}\delta v^{2}/\lambda$ for some fluctuations on scale $\lambda$ (with velocity dispersion $\delta v$ including thermal+drift/bulk components), this is equivalent to assuming $|\delta v^{2}|/\lambda\ll|\omega_{i}v_{\rm drift}|$. This is usually justified by assuming the drift is sub-thermal, and the $\rho_{j}\langle({\bf v}_{j}-\langle{\bf v}_{j}\rangle)({\bf v}_{j}-\langle{\bf v}_{j}\rangle)\rangle$ component of the charge carrier pressure tensors come into thermal equilibrium (are Maxwellian and isotropic with a single temperature, $\delta v\sim v_{T}$) and vary only on macroscopic scales ($\lambda\sim\ell_{\rm macro}$). (3) In the various $\omega_{i}$ terms, assume collisions with any non-neutrals can be neglected ($\omega_{in}\gg\omega_{ij}$ for $j\neq n$), and ignore inelastic and charge-exchange reactions ($\omega^{\rm rec/ion}\ll\omega^{\rm coll}$). (4) Assume the drift ($v_{\rm drift}\equiv{\bf u}_{+}-{\bf u}_{-}$) and slip ($v_{\rm slip}^{\pm}={\bf u}_{\pm}-{\bf u}_{n}$) velocities are vanishingly small compared to the thermal and Alfvén and other wave speeds of all species. This means (a) that the non-isotropic/Reynolds/bulk-flow components of the pressure tensors $\boldsymbol{\Pi}^{\rm aniso}_{j}=\rho_{j}({\bf u}_{j}-{\bf U})({\bf u}_{j}-{\bf U})$ can be neglected, (b) that certain instabilities and anomalous resistivities can be neglected (discussed below), and (c) that cofficients like $\langle\sigma v\rangle_{\rm DF}^{ji}$ (equivalently $\gamma_{ji}$ or $\omega^{\rm coll}_{ji}$) can be taken to have their values for zero drift/slip, which (by definition) are independent of the drift speeds and therefore no longer complicated functions of ${\bf J}$, ${\bf J}_{d}$, ${\bf U}$, and their derivatives. (5) Assume that fluctuations on unresolved (micro/meso) scales are vanishingly small, so all $\delta{\bf X}$ terms in Eq. 7 can be neglected and all coefficients like $\omega_{i}$ are functions only of the mean macro-scale average values $\bar{\bf B}$, $\bar{\bf J}$, etc. Note that this assumption is often implied to be equivalent to the scale hierarchy assumption $\Delta x_{\rm macro}\gg\Delta x_{\rm micro}$, but from the discussion in § 2.1 it is obvious that is not sufficient.
@@
With these assumptions we finally obtain a relatively simple equation for $\bar{\bf B}$:
@@
The $a^{0}_{i}$ notation denotes the value of $a_{i}(\bar{\boldsymbol{\Psi}})$ assuming all state variables have their macroscopic mean values, and dropping terms and simplifying by assuming negligible drift/slip speeds per assumptions (1)-(5).
@@
[…]

### v2

@@ 2.1 Derivation
Figure 1: Illustration of the regimes in density $\bar{n}=\bar{\rho}/m_{p}$ and magnetic field strength $\bar{B}\equiv|\bar{\bf B}|$, in a weakly-ionized gas ($n_{\pm}\ll n$), where different non-ideal MHD terms derived in § 2.1 (Ohmic resistivity $\eta_{O}{\bf J}$, Hall $\eta_{H}{\bf J}\times{\bf b}$, and ambipolar drift $\eta_{A}{\bf J}\times{\bf b}\times{\bf b}$) would each be relatively more important in the induction equation. Quantities like temperature, $v_{T}$, $\omega_{\rm coll}$, and ionization fractions at each $B$, $n$ are estimated as described in § 2.4.2-2.4.3, for a multi-species (electrons, multiple ions, dust grains) system. Left: The case with “classical” non-ideal coefficients (§ 2.2, Eq. 8), which are valid only if the drift velocity between charge carriers $v_{\rm drift}\sim|{\bf J}|/n_{-}|q_{-}|\sim cB/4\pi n_{-}|q_{-}|\ell_{B}$ ($\ell_{B}\sim|{\bf B}|/|\nabla{\bf B}|$) and “slip” velocity between neutrals and carriers (Eq. 9) are much less than the thermal velocities $v_{T}$ ($v_{\rm drift}\ll v_{T}$). This in turn requires that the magnetic gradient scale length $\ell_{B}$ is much larger than some $\ell_{\rm crit}$ (Eq. 13). The multi-species coefficients come from NICIL (§ 2.4.3) and follow the expressions in Appendix A. Right: The case if the implied drift becomes superthermal: $v_{\rm drift}\gtrsim v_{T}$. We use the corrected “effective” coefficients (§ 2.3; Eqs. 15-16), which account for superthermal drift and its effect both on direct particle collision rates and effective collisions via scattering (anomalous resistivity), assuming drift velocities are superthermal (here imposing $v_{\rm drift}=10\,v_{T}$ for illustration), or $\ell_{B}\lesssim\ell_{\rm crit}$ (strong magnetic field gradients), with the effective $\eta_{O}^{\rm an}$ and $v_{T}$ for a multi-species dusty fluid following § 2.4.2, Eqs. 19-20. When the drift becomes superthermal, the effective Ohmic term becomes larger than the Hall term, and increases greatly in relative importance even at low densities. Where it predominates, we label whether the anomalous ($\omega_{\rm an}$) or direct collision ($\omega_{\rm coll}$) term is most important.
We begin with a derivation of the non-ideal MHD equations, following e.g. Cowling (1976); Ichimaru (1978); Nakano & Umebayashi (1986). This has been done before, but this helps motivate our discussion below, particularly regarding which terms can be dropped. Convenient variables used throughout this paper are defined in Table 2.1.
We follow the usual kinetic approach by computing moments of the distribution function of each species, assuming infinitesimally small averaging volumes (but implicitly larger than the interparticle separation and Debye length), with bulk velocities $\ll c$, quasi-neutrality, and collisional/exchange reactions being mass and momentum conserving (e.g. rest mass loss via photons is negligible). For an arbitrary set of species $j$ (which should formally include all isotopes/masses/ionization states/charges) this gives,
@@
Variables Used Throughout
Table 1:
Name Definition $\bf E$}, $\bf J$}, $c$ electric field, current, speed-of-light $\bf B$}, $B$, $\bf b$} magnetic field ${\bf B}\equiv B\,{\bf b}$ $q_{j}$, $m_{j}$, $\rho_{j}$ charge, mass, density of species $j$ $n_{j}$, $T_{j}$, ${\bf u}_{j}$ number density, temperature, bulk velocity of species $j$ $\omega_{ij}^{\rm coll}$, $\gamma_{ij}$ collision rates & rate coefficients between $i$ & $j$ $\tilde{\beta}_{j}$ Hall or mobility parameter of species $j$ $\xi_{j}$, $\Omega_{j}$ mass fraction & gyro frequency of species $j$ $\rho$, $\bf U$} total density and mean velocity (includes all $j$) $\bar{X}$, $\delta{\bf X}$ macroscopic average of $\bf X$}, fluctuations $\delta{\bf X}$ $\tilde{\bf J}_{A}$ normalized Ampere’s current $\tilde{\bf J}_{A}\equiv\nabla\times\bar{\bf B}$ $\eta_{O,\,H,\,A}$ Ohmic, Hall, ambipolar diffusivities $\eta_{O,\,H,\,A}^{0}$ diffusivities in the standard approximation $\eta_{O}^{\rm an}$, $\omega^{\rm an}$ diffusivity, effective collisions from anomalous resistivity $\eta_{O,\,H,\,A}^{\rm eff}$ approximate $\eta$ with anomalous & finite-drift effects ${\bf v}_{\rm drift}$ drift velocity between charge-carriers ${\bf v}^{\pm}_{\rm slip}$ slip speed between neutrals and charged ($\pm$) species $v_{T}$ characteristic slow wavespeed (e.g. thermal speed) $\ell_{B}$ magnetic gradient scale-length $\ell_{\rm crit}^{\rm drift}$, $\ell_{\rm crit}^{\rm slip}$ value of $\ell_{B}$ below which $v_{\rm drift/slip}>v_{T}$
Now consider a three-component system with a dominant set of neutrals, positive and negative charge carriers (labeled $n$, ++, and -, respectively). We can repeat this derivation for an arbitrary set of species, and we do so in Appendix A to validate our expressions. However in the general multi-species case many of our key results must be calculated numerically, while we can derive closed-form analytic expressions for the three-component case which inform our intuition and capture the salient dimensional scalings.
Define an “ambipolar current” ${\bf J}_{d}\equiv-(q_{-}\rho_{c}/m_{-})\,({\bf u}_{+}-{\bf u}_{n})$, where $\rho_{c}\equiv\rho_{+}+\rho_{-}$. We can then express the three velocities ${\bf u}_{n,+,-}$ in terms of $\bf U$}, $\bf J$}, and ${\bf J}_{d}$: ${\bf u}_{n}={\bf U}-(m_{-}/q_{-}\rho)\,({\bf J}-{\bf J}_{d})$, ${\bf u}_{+}={\bf U}-(m_{-}/q_{-}\rho)\,[{\bf J}+(\rho_{n}/\rho_{c}){\bf J}_{d}]$, ${\bf u}_{-}={\bf U}-(m_{-}/q_{-}\rho)\,[-\{(\rho_{+}+\rho_{n})/\rho_{-}\}{\bf J}+(\rho_{n}/\rho_{c}){\bf J}_{d}]$ (Cowling, 1976; Pinto et al., 2008), and rewrite our three Eq. 1 for the evolution of ${\bf u}_{n,+,-}$ with an equivalent set for $\bf U$}, $\bf J$}, ${\bf J}_{d}$ via substitution. With Eq. 2 for $\bf U$}, we have:
@@
${\bf G}_{0}$ $\equiv\frac{|q_{-}|}{m_{-}}[{\xi_{n}\nabla\cdot\boldsymbol{\Pi}-\nabla\cdot\boldsymbol{\Pi}_{n}+\xi_{n}({\bf a}_{e,-}\rho_{-}+{\bf a}_{e,+}\rho_{+}-{\bf a}_{e,n}\rho_{c})}]\,$,
@@
$\qquad+{{\bf D}_{t0}}+{\bf G}_{2}-\frac{|q_{-}|}{m_{-}}\nabla\cdot(\boldsymbol{\Pi}_{-}+\epsilon\boldsymbol{\Pi}_{+}){\Bigr]}\ ,$
@@
Note that $\omega_{O}$ is similar to an effective collision frequency $\omega^{\rm coll}$; $\omega_{H}$ is similar to the gyrofrequency $\sim\Omega_{-}$, and $\omega_{A}$ scales as $\sim\Omega_{-}\Omega_{+}/\omega^{\rm coll}$.
Eq. 5 (with $\partial_{t}{\bf B}=-c\nabla\times{\bf E}$) formally gives us an expression for the evolution of $\bf B$}, but this is clearly not solveable exactly without kinetic methods that can resolve the plasma “micro” scales (e.g. gyro radii, collisional mean-free paths, etc.) and separately predict terms like the stress tensors $\boldsymbol{\Pi}_{j}$, time derivatives $D_{t}{\bf J}$, $D_{t}{\bf J}_{d}$, battery terms, etc. Even if one adopted some effective closure relations for these terms in Eqs. 1-5 to attempt to explicitly integrate a “three fluid” method, the equations could, in many cases of interest, develop micro-scale (kinetic) or meso-scale (small-scale fluid) fluctuations in $\bf B$} or $\bf J$}. As we are interested in the dynamics of large (“macro”) scales in the system, these could also contribute importantly to the large-scale dynamics. Thus, a crucial set of assumptions made in order to justify dropping certain terms in Eq. 5 and therefore render it integrable on “macro” scales $\ell\sim\Delta x_{\rm macro}$, $T\sim\Delta t_{\rm macro}$ is to assume some scale separation of macro/meso/micro scales, and then to assume that fluctuations on the micro and meso-scales are negligible for the macro-scale quantities.
Mathematically, one can represent this with some Reynolds expansion. Heuristically, define the large-scale average of some quantity $\bf X$} as $\bar{\bf X}\equiv\langle{\bf X}\rangle_{\mathcal{V}}=(\int_{\mathcal{V}}{\bf X}\,d^{3}{\bf x}\,dt)/(\int_{\mathcal{V}}d^{3}{\bf x}\,dt)$, integrating over some macro-scale hyperspace domain $\mathcal{V}$ which is large compared to micro/meso scales, but still small compared to global length scales of the problem. Then define ${\bf X}\equiv\bar{\bf X}+\delta{\bf X}$ where $\delta{\bf X}({\bf x},\,t)$ represents mesoscale fluctuations (and $\langle{\delta{\bf X}}\rangle_{\mathcal{V}}=0$). One can then simply insert these definitions to expand the induction equation $\partial_{t}{\bf B}=-c\nabla\times{\bf E}$ and obtain:
$\partial_{t}\bar{\bf B}$ $\approx-c\left[\nabla\times\bar{\bf E}+\langle\nabla\times\delta{\bf E}\rangle_{\mathcal{V}}\right]$ (7)
@@
$+\langle\nabla\times\left[\delta{\bf U}\times\bar{\bf B}+\bar{\bf U}\times\delta{\bf B}+...\right]\rangle_{\mathcal{V}}$
[…]
@@ 2.2 The “Standard” Non-Ideal MHD Approximations
To arrive at the standard formulation of one/two-fluid non-ideal MHD, we simplify Eq. 5 or 7 by making the following assumptions (see Ichimaru, 1978; Norman & Heyvaerts, 1985; Nakano & Umebayashi, 1986; Wardle & Ng, 1999; Tassis & Mouschovias, 2005, 2007; Kunz & Mouschovias, 2009). (1) Assume (given dimensional arguments since velocities are non-relativistic: $|{\bf U}|,\,c_{s},\,v_{\rm drift},\,v_{A}\ll c$) that displacement currents can be neglected so ${\bf J}\rightarrow{\bf J}_{A}\equiv(c/4\pi)\nabla\times{\bf B}$. (2) Drop all terms from the second-line of Eq. 5. (a) Assume temporal/advection frequencies $D_{t}$ are much smaller than the collision+gyro frequencies $\omega_{i}$ ($|D_{t}|\ll\omega_{i}$), usually justified by assuming $|D_{t}|\sim 1/\Delta t_{\rm macro}$ varies only on “macroscopic” timescales because all currents come into local steady-state ($D_{t}\bar{\bf X}\rightarrow 0$) rapidly. (b) Assume that the battery terms are negligible, i.e. external accelerations are negligible compared to the typical non-ideal terms, $|{\bf a}_{\rm ext}|\ll|\omega_{i}v_{\rm drift}|$, ionization fractions are small, and the difference between external acceleration/pressure forces on neutrals and charge carriers is small. (c) Assume the stresses are similarly negligible. Noting $\nabla\cdot(\boldsymbol{\Pi}_{-}+\epsilon\boldsymbol{\Pi}_{+})\sim\rho_{-}\delta v^{2}/\lambda$ for some fluctuations on scale $\lambda$ (with velocity dispersion $\delta v$ including thermal+drift/bulk components), this is equivalent to assuming $|\delta v^{2}|/\lambda\ll|\omega_{i}v_{\rm drift}|$. This is usually justified by assuming the drift is sub-thermal, and the $\rho_{j}\langle({\bf v}_{j}-\langle{\bf v}_{j}\rangle)({\bf v}_{j}-\langle{\bf v}_{j}\rangle)\rangle$ component of the charge carrier pressure tensors come into thermal equilibrium (are Maxwellian and isotropic with a single temperature, $\delta v\sim v_{T}$) and vary only on macroscopic scales ($\lambda\sim\ell_{\rm macro}$). (3) In the various $\omega_{i}$ terms, assume collisions with any non-neutrals can be neglected ($\omega_{in}\gg\omega_{ij}$ for $j\neq n$), and ignore inelastic and charge-exchange reactions ($\omega^{\rm rec/ion}\ll\omega^{\rm coll}$). (4) Assume the drift ($v_{\rm drift}\equiv{\bf u}_{+}-{\bf u}_{-}$) and slip ($v_{\rm slip}^{\pm}={\bf u}_{\pm}-{\bf u}_{n}$) velocities are vanishingly small compared to the thermal and Alfvén and other wave speeds of all species. This means (a) that the non-isotropic/Reynolds/bulk-flow components of the pressure tensors $\boldsymbol{\Pi}^{\rm aniso}_{j}=\rho_{j}({\bf u}_{j}-{\bf U})({\bf u}_{j}-{\bf U})$ can be neglected, (b) that certain instabilities and anomalous resistivities can be neglected (discussed below), and (c) that cofficients like $\langle\sigma v\rangle_{\rm DF}^{ji}$ (equivalently $\gamma_{ji}$ or $\omega^{\rm coll}_{ji}$) can be taken to have their values for zero drift/slip, which (by definition) are independent of the drift speeds and therefore no longer complicated functions of $\bf J$}, ${\bf J}_{d}$, $\bf U$}, and their derivatives. (5) Assume that fluctuations on unresolved (micro/meso) scales are vanishingly small, so all $\delta{\bf X}$” terms in Eq. 7 can be neglected and all coefficients like $\omega_{i}$ are functions only of the mean macro-scale average values $\bar{\bf B}$, $\bar{\bf J}$, etc. Thus for all variables ${\bf X}\rightarrow\bar{\bf X}$. Note that this assumption is often implied to be equivalent to the scale hierarchy assumption $\Delta x_{\rm macro}\gg\Delta x_{\rm micro}$, but from the discussion in § 2.1 it is obvious that is not sufficient.
@@
With these assumptions, applied to Eq. 5 to obtain ${\bf E}\approx\bar{\bf E}$ and $\partial_{t}\bar{\bf B}\approx-c\nabla\times\bar{\bf E}$ , we finally obtain a relatively simple equation for $\bar{\bf B}$:
@@
The $a^{0}_{i}$ notation denotes the value of $a_{i}(\bar{\boldsymbol{\Psi}})$ (Eq. 6) assuming all state variables have their macroscopic mean values, and dropping terms and simplifying by assuming negligible drift/slip speeds per assumptions (1)-(5).
@@
where the latter dimensional scalings come from taking the mostly-neutral limit for $\omega_{O}^{0}/\Omega_{1}^{0}\approx\omega_{\rm coll}^{O,\,0}/\bar{\Omega}_{-}$ with $\omega_{\rm coll}^{O,\,0}=\omega_{+n}^{0}\omega_{-n}^{0}/\omega_{\rm coll}^{A,\,0}$ and $\omega_{\rm coll}^{A,\,0}=\omega_{+n}^{0}+\epsilon\omega_{-n}^{0}$; $\omega_{H}^{0}/\Omega_{1}^{0}\approx(\omega_{+n}^{0}-\epsilon\omega_{-n}^{0})/\omega_{\rm coll}^{A,\,0}\sim\pm 1$; and $\omega_{A}^{0}/\Omega_{1}^{0}\approx\bar{\Omega}_{+}/\omega_{\rm coll}^{A,\,0}$.